# BatchLLM: Optimizing Large Batched LLM Inference with Global Prefix Sharing and Throughput-oriented Token Batching

### v1

@@ 4.2. Explicit Global KV Reuse Identification
As discussed in Sec.3.1, different prompts may have multiple levels of prefixes. The multi-level prefixes can lead to system challenges and overhead of token batching and fused Attention kernel. For the token batching, it requires to manage the dependencies of the different levels of prefixes. For the Attention kernel, it will introduce more GPU kernels of the multi-level reuse. Instead, BatchLLM compresses the multi-level prefixes into a single level prefix, to avoid the above challenges. This is based on the insight that the length of the prefixes are usually dominated by a single level due to the stronger reasoning ability of the newer LLMs and the task-specific fine-tuning (e.g., the prefix corresponding to the web document body can be much longer than other levels of the prefix, as discussed in Sec.3.1).
Figure 4. The preprocessing to maximize the first level prefix reusing. It converts from the initial prefix tree (a) to the final prefix tree (c). Each cycle represents a node containing a number of tokens. The number in the cycle is the token number of the node. It iterates bottom-up to maximize the first level reusing recursively until reaching the root node, which is a dummy node.
As discussed in Sec.3.3, BatchLLM explicitly identifies the common prefix among the prompts within the large batch. It makes use of the compact prefix tree [^5] The compact prefix tree, also called Radix tree, is the prefix tree in which each node that is the only child is merged with its parent. to identify and represent the common prefix between prompts. However, directly using the first level prefix of the basic prefix tree as the common prefix can lead to suboptimal prefix reusing. Take the pesudo prompts in Fig.4d as an example. The 6 prompts are organized with the compact prefix tree in Fig.4a, where each node represents several consecutive tokens of the prompt (the number in the node means the number of tokens) and the tokens in a node is shared by its children. The right child of the root node (i.e., token ’b’) is shared by 5 prompts in total (prompt 2 to 6) according to Fig.4a, thus reusing the first level prefix can save 4 tokens’ KV computation. However, it is easy to infer that sharing the first 8 tokens between prompt 4 and 5 can save 8 tokens’ KV computation. Thus it requires to refactor the naive prefix tree to maximize the achieved first-level token reusing.
@@
1$\triangleright$ Note the 1st-level prefixes of a node are its children’s tokens
@@
We have compared the token saving ratio between the original multi-level prefix and the enlarged single prefix. The saving ratio is calculated according to Eq.2. For a dataset of the snippet generation task, the saving ratio of the original multi-level prefixes is about 56%, and the ratio is about 55% for the first-level prefix after enlarging (Sec.6.2.2). It means the enlarged first-level prefix only has 1% loss of the token reuse compared to the multi-level prefix for this typical workload, with the advantage of reducing the complexity and the overhead of the token batching and Attention kernel optimization.
@@ 4.3.1. Prefix-sharing Group Based Scheduling
As described in Sec.4.2, BatchLLM identifies the common prefix explicitly ahead-of-time. To simplify the KV reuse and reduce the lifetime of the common prefix in the memory, BatchLLM schedules the requests sharing the same prefix all together. Specifically, the procedure described in Sec.4.2 will generate the prefix-sharing groups and BatchLLM will schedule the requests at the granularity of prefix-sharing group rather than each single request. A prefix-sharing group corresponds to the collection of requests sharing the same prefix. In this way, the requests sharing the same prefix can start at the same time and the memory of the common prefix can be released as soon as the group is completed. This differs from the LRU cache and reference counter based prefix management (Kwon et al., 2023; Zheng et al., 2023) and can make sure all the common prefix can be reused the best and the memory will not be retained unnecessarily.
To realize the prefix-sharing group based scheduling, BatchLLM maintains three queues of the to-be-batched tokens: common prefix queue, distinct prompt queue, and decoding queue. At each iteration, BatchLLM will form the token-batch by fetching the tokens from the three queues in the order of decoding queue, distinct prompt queue, and common prefix queue (one constraint is that the a common prefix should be fetched and processed before its corresponding distinct prompts). On the one hand, fetching the decoding tokens before prefills can help to better mix the two types of tokens together in the token-batch (Holmes et al., 2024). On the other hand, fetching the decoding and distinct prompt tokens before the prefix tokens can make sure the active requests (i.e., the request that has been partially processed, either partial prefill chunks has been processed or already in decoding phase) can be scheduled before the inactive requests, which helps to finish the active requests earlier and thus release their memory earlier.
@@ 4.3.2. Request Groups Reordering
In Eq.3, $L_{prefill}$ is the number of prefill tokens and $L_{decode}$ is the length of the output tokens. The requests with larger $R$ will be scheduled earlier. This helps to issue the requests with long decoding steps earlier and make the latter prompts’s length larger and thus can be better mixed with the previous decoding tokens to enlarge the token-batch size (example in Fig.1). The challenge is that the exact number of output tokens is not known before finishing the execution. We profiled the dataset of several typical information processing tasks (snippet generation, offline ranking, ads, etc.) and observed that the distribution of output length is relatively concentrated for a task. We thus assume the output length is constant and normalize $L_{prefill}$ into 1 in Eq.3 for all requests uniformly. As we schedule the requests in the unit of prefix-sharing group, the ratio of the group (after normalizing the length of output) is
@@ 4.3.3. Memory-centric Token Batching to Saturate GPU
Specifically, BatchLLM predefines a size as the upperbound of the KV memory size ($M_{threshold}$). When deciding the token-batching at each iteration, it will only check whether adding a prefill chunk will exceed $M_{threshold}$ or not, without considering the number of requests. The prefill chunk will be added into the token-batch if the remaining GPU memory capacity allows it. Algo.2 describes the detailed procedure. Note that we also use a fixed number to limit the per-batch token number for token-batching decision (noted as the chunk size $S_{chunk}$), which aligns with that in vLLM and helps to distribute the token numbers in different iterations more even.
Algorithm 2 Simplified memory-centric token batching.
1:Chunk size $S_{chunk}$, KV memory threshold $M_{threshold}$.
2:$RaggedBatch$ is the current token-batch containing the active decoding tokens
3:$M_{current}$ is the current usage of KV memory
4:while $|RaggedBatch|<S_{chunk}$ and $M_{current}<M_{threshold}$ do
5:     if prefill request $p$ available then
6:         Add $p$ to $RaggedBatch$
7:         $M_{current}\leftarrow M_{current}+M_{p}$
8:     else
9:         Break
10:return $RaggedBatch$
4.4. Prefix-shared Attention Optimization
As described in Sec.3.2, the existing works use separate kernels to compute the partial Attention on the prefix and the distinct KV context. This will lead to the tail effect of each kernel and the launch overhead of these kernels. BatchLLM horizontally fuses the two partial Attention calculation into one kernel. Different parts of the computation are performed in different thread blocks. Note that the problem shape of the two parts are different. BatchLLM applies different tiling configuration for the two parts to achieve the best performance. It uses the common auto-tuning method to find the best configuration.
The horizontal fusion itself is not novel, but can effectively help to increase the performance of the prefix-shared Attention.
@@ 5. Implementation
We integrate the token batching optimization (Sec.4.3) in vLLM (Kwon et al., 2023) v0.5.4. Specifically, we customize the vLLM to accept the list of prefix-sharing group tuples generated by the preprocessing script, and implement the group-wised scheduling and token batching logic upon the vLLM token batching function.
Finally, we integrate the prefix-shared Attention optimization (Sec.4.4) as the Attention kernel backend of the customized vLLM. We implement the Attention in BatchLLM through OpenAI Triton language (Tillet et al., 2019). The limitation of the Triton based implementation is that, the performance of Triton version of FlashAttention is still worse than that of CUDA version (Sec.6.3.3). However, the advantage is that Triton supports multiple platforms (both NVIDIA and AMD GPUs) so that BatchLLM implementation can run on both NVIDIA and AMD GPUs. We use the builtin autotuner of Triton to tune the tiling size of the Attention implementation in BatchLLM. Besides that, some ahead-of-time compilation methods are applied to reduce the launching overhead, including warm-up process on NVIDIA GPUs, and AOTriton [^7] https://github.com/ROCm/aotriton on AMD GPUs.
@@ 6.1. Setup
Baseline Specification. We compare BatchLLM with vLLM, the state-of-the-art LLM inference engine with prefix sharing, for the end-to-end comparison to demonstrate the efficacy (Sec.6.2). We integrate all our design in vLLM v0.5.4 and the baseline for comparison is also vLLM v0.5.4 (except as noted). We use FlashAttentionn (Dao et al., 2022) (based on v2.6.1) as the Attention backend888 We do not use Flashinfer (Ye et al., 2024a) as the backend as we observe it will result in incorrect results of the end-to-end output in vLLM v0.5.4. Note that the single kernel result of Flashinfer is correct, so that there may be some engineer problem of its integration to vLLM v0.5.4. of the vLLM baseline in NVIDIA GPU environments.
We enable the prefix-caching and chunked-prefill in vLLM baseline for the end-to-end comparison in Sec.6.2. The prefix-caching uses the implicit LRU-based caching policy. We use 2048 as the token-batch size when enabling the chunked-prefill of vLLM baseline, which is a recommended size to achieve its better throughput according to the vLLM community. Note that the prefix-caching and chunked-prefill are still not supported to be enabled together in v0.5.4. To make a fair comparison, we optimized the code of vLLM 0.5.4 baseline to make it able to enable the two techniques together, which makes the vLLM baseline have better performance in Sec.6.2.
For the kernel comparison in Sec.6.3.3, we compare BatchLLM with Cascade-Inference (Ye et al., 2024a), which is the SOTA implementation of the prefix-shared Attention on NVIDIA platform. It is worth noting that Cascade-Inference is similar to our design on kernel-levels. One difference between Cascade-Inference and BatchLLM is that BatchLLM applies horizontal fused kernel as described in 4.4. Besides, Cascade-Inference cannot be applied on the non-NVIDIA backend as it is developed with CUDA, while BatchLLM is based on OpenAI Triton and can support both NVIDIA and AMD GPUs.
All the other packages like PyTorch and Huggingface transformers are dependent to the docker file of vLLM v0.5.4. The OpenAI Triton version is 3.0.0.
@@
Models. We evaluate BatchLLM and the baselines with Qwen2-7B (Yang et al., 2024), Mistral-7B (Jiang et al., 2023) and Llama3-8B (AI@Meta, 2024) models. These are among the most popular models with the size affordable for the information processing and management systems dealing with massive inputs.
Dataset. We evaluate BatchLLM and the baselines with both three microbenchmarks with different data distributions and two typical industry tasks.
As for the microbenchmarks evaluation, we use three different settings of the length of common prefix and distinct prompt. The length of prefix/distinct are 2000/200, 1000/1000 and 200/2000 respectively. Note that the lengths are typical for the industry workloads. We evaluate the three microbenchmarks under different share-degrees. The share-degree means the number of distinct prompts sharing the same prefix. Specifically, we conduct the evaluation of the three microbenchmarks with sharing degree 4 and 16, respectively. The microbenchmark evaluation is in Sec.6.2.1.
As for the industry tasks, we conduct the evaluation on two typical scenarios in web search engines. The one is that, given a document (as the shared prefix), the task is to extract the corresponding information according to the user query (e.g., snippet generation task). The other is that, given a user query (as the shared prefix), the task is to rank a list of documents according to some metrics (e.g., offline ranking). These two scenarios have different data distribution so that it can better demonstrate the efficacy of BatchLLM. Specifically, the common prefix of the former task is much longer than the distinct prompt, and the latter task is exactly the opposite. We construct two datasets synthesized according to two of the above scenarios in industry, and make the case study in 6.2.2 and 6.2.3.
@@ 6.2.1. Microbenchmark Evaluation
(a) Llama3-8B Throughput, on A100.
(b) Qwen2-7B Throughput, on A100.
(c) Mistral-7B Throughput, on A100.
(d) Qwen2-7B Throughput, on MI200.
(e) Mistral-7B Throughput, on MI200.
Figure 5. Microbenchmark evaluation of different models and hardware enviromments. The setting m/n (like 200/2000) indicates the length of shared prefix and non-shared context. sd means shared-degree, showing that how many requests in one prefix-sharing group with the same shared prefix. In the legend, the setting with ’+ p’ means prefix-caching enabled, while ’+ c’ means chunked-prefill enabled.
Fig.5 shows the end-to-end throughput comparison between BatchLLM and the baselines. In these benchmark experiments, we compare BatchLLM with vLLM of different configurations, under the FlashAttention backend (Dao et al., 2022), and with or without prefix-caching and chunked-prefill enabled.
The results show that BatchLLM outperforms the vLLM baseline under different configurations, for different models on different GPU platforms with different input data distributions. Specifically, when comparing with vLLM + prefix-caching + chunked-prefill, which performs the best across all vLLM’s configurations, BatchLLM can get 1.1-1.5$\times$ speedup on A100 and up to 2$\times$ speedup on MI200. It also shows that the speedup becomes more significant when the portion of the common prefix becomes larger. This is because the reuse ratio becomes larger and BatchLLM can save more computation and memory for this case. The 200/2000 dataset has little room of KV reusing and thus all the optimizers do not show significant performance improvement.
The performance improvement of BatchLLM comes from the better KV reuse with the explicit prefix identification, the token-batching to better mix the decoding with prefill chunks together, and the better Attention kernels, which we will analyze through the breakdown in Sec.6.3.
@@
Settings Case 1 (CUDA) Case 2 (ROCm)
vLLM 7.13 4.34
+ prefix-caching 9.9 4.41
+ chunked-prefill 7.08 4.56
+ prefix-caching + chunked-prefill 9.83 5.88
BatchLLM 12.85 8.65
6.2.2. Case Study 1: Long Common Prefix with Short Distinct Prompt.

### v2

@@ 4.2. Explicit Global KV Reuse Identification
As discussed in Sec.3.1, different prompts may have multiple levels of prefixes. The multi-level prefixes can lead to system challenges and overhead of token batching and fused Attention kernel. For the token batching, it requires to manage the dependencies of the different levels of prefixes. For the Attention kernel, it will introduce more GPU kernels of the multi-level reuse. Instead, BatchLLM compresses the multi-level prefixes into a single level prefix to avoid the above challenges. This is based on the insight that the length of the prefixes are usually dominated by a single level due to the stronger reasoning ability of the newer LLMs and the task-specific fine-tuning (e.g., the prefix corresponding to the web document body can be much longer than other levels of the prefix, as discussed in Sec.3.1).
As discussed in Sec.3.3, BatchLLM explicitly identifies the common prefix among the prompts within the large batch. It makes use of the compact prefix tree [^5] The compact prefix tree, also called Radix tree, is the prefix tree in which each node that is the only child is merged with its parent. to identify and represent the common prefix between prompts. However, directly using the first level prefix of the basic prefix tree as the common prefix can lead to suboptimal prefix reusing. Take the pesudo prompts in Fig.4d as an example. The 6 prompts are organized with the compact prefix tree in Fig.4a, where each node represents several consecutive tokens of the prompt and the tokens in a node is shared by its children. The right child of the root node (i.e., token ’b’) is shared by 5 prompts in total (prompt 2 to 6) according to Fig.4a, thus reusing the first level prefix can save 4 tokens’ KV computation. However, it is easy to infer that sharing the first 8 tokens between prompt 4 and 5 can save 8 tokens’ KV computation. Thus it requires to refactor the naive prefix tree to maximize the achieved first-level token reusing.
@@
1$\triangleright$ The 1st-level prefixes of a node are its children’s tokens
@@
We have compared the token saving ratio between the original multi-level prefix and the enlarged single prefix. The saving ratio is calculated according to Eq.2. For a dataset of the snippet generation task, the saving ratio of the original multi-level prefixes is about 56%, and the ratio is about 55% for the first-level prefix after enlarging (Sec.6.2.2). It means the enlarged first-level prefix only has 1% loss of the token reuse compared to the multi-level prefix for this typical workload, with the advantage of reducing the complexity and the overhead of the token batching and Attention kernel optimization. Besides, the DP algorithm only introduces less than 0.01% overhead in our evaluation (Sec.6.3.4).
Figure 4. The preprocessing to maximize the first level prefix reusing. It converts from the initial prefix tree (a) to the final prefix tree (c). Each cycle represents a node containing a number of tokens. The number in the cycle is the token number of the node. It iterates bottom-up to maximize the first level reusing recursively until reaching the root node.
@@ 4.3.1. Prefix-sharing Group Based Scheduling
As described in Sec.4.2, BatchLLM identifies the common prefix explicitly ahead-of-time. To simplify the KV reuse and reduce the lifetime of the common prefix in the memory, BatchLLM schedules the requests sharing the same prefix all together. Specifically, the procedure described in Sec.4.2 will generate the prefix-sharing groups and BatchLLM will schedule the requests at the granularity of prefix-sharing group rather than each single request. A prefix-sharing group corresponds to the collection of requests sharing the same prefix. In this way, the requests sharing the same prefix can start at the same time and the memory of the common prefix can be released as soon as the group is completed. This differs from the LRU cache based prefix management (Kwon et al., 2023; Zheng et al., 2023) and can make sure all the common prefix can be reused the best and the memory will not be retained unnecessarily.
To realize the prefix-sharing group based scheduling, BatchLLM maintains three queues of the to-be-batched tokens: common prefix queue, distinct prompt queue, and decoding queue. This method differs from the existing methods that use the same queue for all prefill tokens. At each iteration, BatchLLM will form the token-batch by fetching the tokens from the three queues in the order of decoding queue, distinct prompt queue, and common prefix queue (one constraint is that the a common prefix should be fetched and processed before its corresponding distinct prompts). On the one hand, fetching the decoding tokens before prefills can help to better mix the two types of tokens together in the token-batch (Holmes et al., 2024). On the other hand, fetching the decoding and distinct prompt tokens before the prefix tokens can make sure the active requests (i.e., the request that has been partially processed, either partial prefill chunks has been processed or already in decoding phase) can be scheduled before the inactive requests, which helps to finish the active requests earlier and thus release their memory earlier.
@@ 4.3.2. Request Groups Reordering
In Eq.3, $L_{prefill}$ is the number of prefill tokens and $L_{decode}$ is the length of the output tokens. The requests with larger $R$ will be scheduled earlier. This helps to issue the requests with long decoding steps earlier and make the latter prompts’ length larger and thus can be better mixed with the previous decoding tokens to enlarge the token-batch size (example in Fig.1). The challenge is that the exact number of output tokens is not known before finishing the execution. We profiled the dataset of several typical information processing tasks (snippet generation, offline ranking, ads, etc.) and observed that the distribution of output length is relatively concentrated for a task. We thus assume the output length is constant and normalize $L_{prefill}$ into 1 in Eq.3 for all requests uniformly. As we schedule the requests in the unit of prefix-sharing group, the ratio of the group (after normalizing the length of output) becomes:
@@ 4.3.3. Memory-centric Token Batching to Saturate GPU
Specifically, BatchLLM predefines a size as the upperbound of the KV memory size ($M_{threshold}$). When deciding the token-batching at each iteration, it will only check whether adding a prefill chunk will exceed $M_{threshold}$ or not, without considering the number of requests. The prefill chunk will be added into the token-batch if the remaining GPU memory capacity allows it. Note that we also use a fixed number to limit the per-batch token number for token-batching decision (noted as the chunk size $S_{chunk}$), which aligns with that in vLLM and helps to distribute the token numbers in different iterations more even.
@@ 5. Implementation
Horizontal Fusion of Prefix-shared Attention. As described in Sec.3.2, the existing works use separate kernels to compute the partial Attention on the prefix and the distinct KV context. This will lead to the tail effect of each kernel and the launch overhead of these kernels. BatchLLM horizontally fuses the two partial Attention calculation into one kernel. Different parts of the computation are performed in different thread blocks. Note that the problem shape of the two parts are different. BatchLLM applies different tiling configuration for the two parts to achieve the best performance. It uses the common auto-tuning method to find the best configuration. We implement the Attention in BatchLLM through OpenAI Triton language (Tillet et al., 2019). The limitation of the Triton based implementation is that, the performance of Triton version of FlashAttention is still worse than that of CUDA version (Sec.6.3.3). However, the advantage is that Triton supports multiple platforms (both NVIDIA and AMD GPUs) so that BatchLLM implementation can run on both NVIDIA and AMD GPUs. We use the builtin autotuner of Triton to tune the tiling size of the Attention implementation in BatchLLM. Besides that, some ahead-of-time compilation methods are applied to reduce the launching overhead, including warm-up process on NVIDIA GPUs, and AOTriton [^7] https://github.com/ROCm/aotriton on AMD GPUs.
@@
We integrate the token batching optimization (Sec.4.3) in vLLM (Kwon et al., 2023). Specifically, we customize the vLLM to accept the list of prefix-sharing group tuples generated by the preprocessing script, and implement the group-wised scheduling and token batching logic upon the vLLM token batching function.
(a) Llama 3.1 8B throughput on A100, request-count = 6400
(b) Qwen 2.5 14B throughput on A100, request-count = 6400
(c) Llama 3.1 70B throughput, on A100, request-count = 800
(d) Llama 3.1 8B throughput, on MI200, request-count = 6400
(e) Qwen 2.5 14B throughput, on MI200, request-count = 6400
Figure 5. Microbenchmark evaluation. The setting m/n (like 2000/200) indicates the length of shared prefix/non-shared context, sd means sharing degree. The vLLM setting with ’+ p’ (’+ c’) means prefix-caching (chunked-prefill) enabled.
@@ 6.1. Setup
Baseline Specification. We compare BatchLLM with vLLM and SGLang, two state-of-the-art LLM inference engines with prefix sharing, for the end-to-end comparison to demonstrate the efficacy (Sec.6.2). We integrate all our design in vLLM v0.6.4 and the baseline for comparison is also vLLM v0.6.4 (except as noted). We use FlashAttention (Dao et al., 2022) as the Attention backend888 We also tested FlashInfer (Ye et al., 2025) as the backend of vLLM v0.6.4, and found that the gap between FlashAttention and FlashInfer was quite small, and the end-to-end results of FlashAttention were slightly better than the other one. of the vLLM baseline in NVIDIA GPU environments. The version of SGLang is v0.4.1.
We enable the prefix-caching and chunked-prefill in vLLM baseline for the end-to-end comparison in Sec.6.2. The prefix-caching uses the implicit LRU-based caching policy. We use 2048 as the token-batch size when enabling the chunked-prefill of vLLM baseline to maximize its throughput.
For the kernel comparison in Sec.6.3.3, we compare BatchLLM with Cascade-Inference (Ye et al., 2024a), which is the SOTA implementation of the prefix-shared Attention on NVIDIA platform. It is worth noting that Cascade-Inference is similar to our design on kernel-levels. One difference between Cascade-Inference and BatchLLM is that BatchLLM applies horizontal fused kernel. Besides, Cascade-Inference cannot be applied on the non-NVIDIA backend as it is developed with CUDA, while BatchLLM is based on OpenAI Triton and can support both NVIDIA and AMD GPUs. All the other packages like PyTorch and Huggingface transformers are dependent to the docker file of vLLM v0.6.4.
@@
Models and Dataset. We evaluate BatchLLM and the baselines with Llama-3.1-8B (AI@Meta, 2024), Qwen-2.5-14B (Qwen et al., 2025) and Llama-3.1-70B (4-bit GPTQ), with a set of microbenchmarks with different data distributions and Qwen-2.5-7B for a typical industry task. As for the microbenchmarks evaluation, we use three different settings of the length of common prefix and distinct prompt. The length of prefix/distinct are 2000/200, 2000/1000 and 1000/1000 respectively, while the length of generated tokens is 100. We evaluate the three microbenchmarks under different share-degrees. The share-degree means the number of distinct prompts sharing the same prefix. Note that all requests are randomly shuffled. Specifically, we conduct the evaluation of the three microbenchmarks with sharing degree 4 and 16, respectively. The microbenchmark evaluation is in Sec.6.2.1. As for the industry tasks, we conduct the evaluation on a typical scenario in web search engines, which is the snippet generation task that generates the snippet of a document for the user query.
@@ 6.2. End to End Comparison
Figure 6. Microbenchmark evaluation of different sharing-degrees and different shared-prefix lengths on A100. Given 6400 requests. This evaluation is based on Llama-3.1-8B-Instruct, while the length of non-shared context is 200, and the generated length is 100.
@@ 6.2.1. Microbenchmark Evaluation
Fig.5 shows the end-to-end throughput comparison between BatchLLM and the baselines with different shared-prefix/non-shared context settings. In these benchmark experiments, we compare BatchLLM with vLLM and SGLang of different configurations, and with or without prefix-caching and chunked-prefill enabled.
The results show that BatchLLM outperforms the vLLM and SGLang baseline under different configurations, for different models on different GPU platforms with different input data distributions. Specifically, when comparing with vLLM + prefix-caching + chunked-prefill, which performs the best across all vLLM’s configurations, BatchLLM can get 1.4-3$\times$ speedup on A100 and 1.8-4.1$\times$ speedup on MI200. BatchLLM also achieves better performance than SGLang, with 1.2-2.5$\times$ speedup on A100 and 1.3-2.8$\times$ speedup on MI200. The acceleration is stable from 7B to 70B models.
The impact of sharing-degree and shared prefix. To show the benefit of BatchLLM, we also compare the throughput of different settings after changing the sharing-degree and the length of shared prefix. Fig.6 shows that the speedup becomes more significant when more requests share the same prefix and the portion of the common prefix becomes larger, which contributes to a bigger reuse ratio, and BatchLLM can save more computation and memory for this case. With lower sharing-degree and shorter common prefix (like sd is 2 and the length of prefix is 2000), the acceleration of BatchLLM is 1.7×1.7$\times$1.7 ×/1.5×1.5$\times$1.5 ×, comparing with the best setting of vLLM and SGLang. When the length of shared prefix is 16000, and the share degree is 16, the best performance of vLLM and SGLang are 0.49 and 0.56, while BatchLLM could achieve 10.8×10.8$\times$10.8 ×/9.5×9.5$\times$9.5 × acceleration.
Comparison between vLLM + request sorting and BatchLLM. We perform a preprocess for the setting of vLLM + chunked-prefill + prefix-cahcing: use python.sorted() to gather all requests with the same prefix together999 Here the length of shared prefix/non-shared context/generated tokens are 2000/200/100, while the sharing degree is 16. The model is Llama-3.1-8B. And the GPU is NVIDIA A100.. This can help vLLM to cache the tokens better. After this preprocess, vLLM’s throughput is boosted from 6 to 13.02, while the throughput of BatchLLM is 18. One reason is that vLLM cannot simplify the attention calculation when there are more than 1 common prefix in the batch every step of its inference, while BatchLLM could handle it. Besides, BatchLLM has a better token-batching strategy than vLLM.
Overall, the performance improvement of BatchLLM comes from the better KV reuse with the explicit prefix identification, the token-batching to better mix the decoding with prefill chunks together, and the better Attention kernels, which we will analyze through the breakdown in Sec.6.3.
@@
Settings Throughtput, A100 Throughput, MI200
vLLM 5.45 2.75
      + c (chunked-prefill) 5.48 2.84
      + p (prefix-caching) 6.2 3.27
      + c + p 6.71 3.36
BatchLLM 8.67(1.30×1.30$\times$1.30 ×) 4.24(1.26×1.26$\times$1.26 ×)
Figure 7. The length distribution of shared prefix/non-shared context in the dataset of the industry workload.
6.2.2. Industry Workload Evaluation